# EgoSonics: Generating Synchronized Audio for Silent Egocentric Videos

### v1

@@ Abstract
We introduce EgoSonics, a method to generate semantically meaningful and synchronized audio tracks conditioned on silent egocentric videos. Generating audio for silent egocentric videos could open new applications in virtual reality, assistive technologies, or for augmenting existing datasets. Existing work has been limited to domains like speech, music, or impact sounds and cannot capture the broad range of audio frequencies found in egocentric videos. EgoSonics addresses these limitations by building on the strengths of latent diffusion models for conditioned audio synthesis. We first encode and process paired audio–video data to make them suitable for generation. The encoded data is then used to train a model that can generate an audio track that captures the semantics of the input video. Our proposed SyncroNet builds on top of ControlNet to provide control signals that enables generation of temporally-synchronized audio. Extensive evaluations show that our model outperforms existing work in audio quality, and in our proposed synchronization evaluation method. Furthermore, we demonstrate downstream applications of our model in improving video summarization.
@@ 2 Related Works
Diffusion Models [20, 44, 38] have demonstrated remarkable efficacy in a multitude of generative tasks, spanning image generation, audio synthesis, and video creation. Among these, latent diffusion models like Stable Diffusion (SD) [38] stand out for their ability to perform both forward and reverse processes within the latent space of data, leading to efficient computation and expedited inference. Notably, SD is engineered to execute inference even on standard personal laptop GPU, making it a practical choice for a wide array of users. Alongside diffusion models, other generative architectures such as Variational Autoencoders (VAEs) [48, 3], Generative Adversarial Networks (GANs), and Autoregressive models have also made significant strides in various creative domains, each with their own strengths and uses.
@@ 3.2 Diffusion Model for Audio Synthesis
In our model, we replace the text prompt from the SD with the video embedding $E_{V}$. The mean of all the image embeddings $\frac{1}{N}\sum_{i=1}^{N}E_{V}^{i}$ is passed through a small MLP to map it to the corresponding $c_{t}$ providing some contextual information to the UNet encoder blocks of the SD. As our video embeddings $E_{V}$ are also generated using CLIP, it makes it easier to reuse existing SD models that already operate with CLIP embeddings. More details about the MLP are given in the supplementary.
@@ 3.3 Time-Aware Audio Control using SyncroNet
where, $Q_{h}$, $K_{h}$, $V_{h}$ represents the Query, Key, and Value matrices derived from $h$.
@@
This would help incorporate contextual information from spatially varying signals in the same feature space, allowing the model to generate quite relevant audio.
@@ 3.4 Training and Inference
For training SyncroNet first, the video frames are encoded using the video encoder. The video embedding is then given to the SyncroNet which generates control signals to guide the SD’s UNet decoder. The UNet decoder generates a $64\times 64$ feature representation, which then goes to the VAE decoder to generate the audio spectrogram. The loss is applied between the generated audio spectrogram $E^{\prime}_{A}$ and the ground truth spectrogram $E_{A}$ as shown in fig. 3. The Syncronet training is carried out out as per the objective mentioned in Eq. 5. Training of Syncronet has been done using the same loss functions as ControlNet [53]. We use DDIM [20] for faster and consistent sampling. Upto 1000 time steps were used in the forward diffusion process, and 20 during the denoising. Training was done using AdamW optimizer with a learning rate of $41e-$ 4 for 50 epochs.
During inference, the given video sampled at 30 fps is given to the CLIP video encoder to obtain an image-like video embedding. Syncronet generates the control signals using this video embedding, which is then given to a DDIM sampler [20] which predicts the conditioned encoded feature map for audio spectrogram. This feature map is decoded using Stable Diffusion’s pretrained decoder to obtain the audio spectrogram synchronized with the input video. The pretrained audio super-resolution (ASR) model is used the enhance the generated audio quality and recover the losses occurred due to rescaling. More details on ASR are given in the supplementary. We scale and invert this enhanced audio spectrogram using Griffin-Lim [33] algorithm to obtain the audio waveform of 10 seconds.
@@ 4 Experiments
Table 1: V2A generation results comparing various methods over 3 metrics (FID, IS, VAAS). We have also compared them against the audio sampling rate, maximum frequency support, spectrogram size used, the length of generated audio, and the number of video frames used for guidance. Our method, using 30 fps, sampled at highest frequency can generate the highest resolution spectrogram for a 10 seconds long audio. It also outperforms the existing V2A on FID, IS, and VAAS.
Method Audio SR Max Freq. Spec. Size Audio Length Video Frames FID $\downarrow$ IS $\uparrow$ VAAS $\uparrow$
Im2Wav $16KHz$ $8KHz$- $5sec$ $1-30fps$ $91.73$ $19.15$ $59.24$
Make-an-Audio $16KHz$ $8KHz$ $80\times 624$ $9-10sec$ $1-4$ $83.72$ $22.14$ $64.57$
Diff-Foley $16KHz$ $8KHz$ $256\times 128$ $8sec$ $4fps$ $76.92$ $37.92$ $78.19$
EgoSonics (ours) $22KHz$ $20KHz$ $512\times 512$ $10sec$ $30fps$ 41.14 54.41 92.77
@@
Table 2: Comparing the effect of using audio generated from different methods on improving video summarization task. It is interesting to note that, even the un-synchronized audios generated from our model have sufficient contextual information to outperform the existing methods.
@@
Im2Wav $0.57$
Make-an-Audio $0.68$
Diff-Foley $0.71$
@@ 4.3 Ablation Study
Unlike previous methods that use an average video embedding, or 1-4 fps for conditioning the audio, we used 30 fps to extract the temporal information from the video for guidance. In this section, we analyzed the impact of using 30 frames on AV synchronization. We compared the performance of SyncroNet at different fps including 4 fps used by the current SOTA Diff-Foley [30]. From Table 3, it can be seen that our method is not only able to generate good quality audio spectrograms, but also able to generate a VAAS score of 85. However, as we increase the fps to 15, the alignment score increases to 89, and then to 92.77 with 30 fps.
@@ 5 Conclusion
In this paper, we build a novel audio-visual framework called EgoSonics. Much like how humans effortlessly connect sound with visual stimuli, EgoSonics has the ability to predict synchronized audio for muted Egocentric videos. Leveraging a streamlined architecture that incorporates generation capabilities of Stable Diffusion, EgoSonics holds promise for various applications. The proposed Syncronet can generate control signals from input video frames to provide pixel-level control to Stable Diffusion. Furthermore, we showcased our method’s superior performance over existing methods and illustrate how this approach can enhance downstream applications like video summarization.
@@
Although, EgoSonics can generate good quality audio for egocentric videos, there are a few limitations to our approach. One of the things we observed in some of our generated samples is the misalignment due to the lack of rich visual information, which might result from occlusions. We are also limited by the amount of data available for training. For example, since there are a very few samples of musical instruments in the Ego4D dataset, our model doesn’t perform very well on such videos. We believe that such challenges can be solved by training our model on a large amount of dataset comprising millions of audio-video pairs.
A. EgoSonics

### v2

@@ Abstract
We introduce EgoSonics, a method to generate semantically meaningful and synchronized audio tracks conditioned on silent egocentric videos. Generating audio for silent egocentric videos could open new applications in virtual reality, assistive technologies, or for augmenting existing datasets. Existing work has been limited to domains like speech, music, or impact sounds and cannot capture the broad range of audio frequencies found in egocentric videos. EgoSonics addresses these limitations by building on the strengths of latent diffusion models for conditioned audio synthesis. We first encode and process paired audio–video data to make them suitable for generation. The encoded data is then used to train a model that can generate an audio track that captures the semantics of the input video. Our proposed SyncroNet builds on top of ControlNet to provide control signals that enables generation of temporally-synchronized audio. Extensive evaluations and a comprehensive user study show that our model outperforms existing work in audio quality, and in our proposed synchronization evaluation method. Furthermore, we demonstrate downstream applications of our model in improving video summarization.
@@ 2 Related Works
Diffusion Models [21, 48, 42] have demonstrated remarkable efficacy in a multitude of generative tasks, spanning image generation, audio synthesis, and video creation. Among these, latent diffusion models like Stable Diffusion (SD) [42] stand out for their ability to perform both forward and reverse processes within the latent space of data, leading to efficient computation and expedited inference. Notably, SD is engineered to execute inference even on standard personal laptop GPU, making it a practical choice for a wide array of users. Alongside diffusion models, other generative architectures such as Variational Autoencoders (VAEs) [54, 3], Generative Adversarial Networks (GANs) [17, 25, 39, 40, 53], and Autoregressive models have also made significant strides in various domains, each with their own strengths and uses.
@@ 3.2 Diffusion Model for Audio Synthesis
We found Stable Diffusion’s VAE to be very effective in reconstruction and generating spectrograms. The same observation was made in [32]. In our model, we replace the text prompt from the SD with the video embedding $E_{V}$. The mean of all the image embeddings $\frac{1}{N}\sum_{i=1}^{N}E_{V}^{i}$ is passed through a small MLP to map it to the corresponding $c_{t}$ providing some contextual information to the UNet encoder blocks of the SD. As our video embeddings $E_{V}$ are also generated using CLIP, it makes it easier to reuse existing SD models that already operate with CLIP embeddings. More details about the MLP are given in the supplementary.
@@ 3.3 Time-Aware Audio Control using SyncroNet
where, $Q_{h}$, $K_{h}$, $V_{h}$ represents the Query, Key, and Value matrices derived from $h$. This would help incorporate contextual information from spatially varying signals in the same feature space, allowing the model to generate quite relevant audio.
@@ 3.4 Training and Inference
For training SyncroNet first, the video frames are encoded using the video encoder. The video embedding is then given to the SyncroNet which generates control signals to guide the SD’s UNet decoder. The UNet decoder generates a $64\times 64$ feature representation, which then goes to the VAE decoder to generate the audio spectrogram. The loss is applied between the generated audio spectrogram $E^{\prime}_{A}$ and the ground truth spectrogram $E_{A}$ as shown in fig. 3. The Syncronet training is carried out out as per the objective mentioned in Eq. 5. Training has been done using the same loss functions . We use DDIM [21] for faster and consistent sampling. Upto 1000 time steps were used in the forward diffusion process, and 20 during the denoising. Training was done using AdamW optimizer with a learning rate of $41e-$ 4 for 50 epochs. Once, the model is trained, we can sample new objects from the learned
During inference, the given video sampled at 30 fps is given to the CLIP video encoder to obtain an image-like video embedding. Syncronet generates the control signals using this video embedding, which is then given to a DDIM sampler [21] which predicts the conditioned encoded feature map for audio spectrogram. This feature map is decoded using Stable Diffusion’s pretrained decoder to obtain the audio spectrogram synchronized with the input video. The pretrained audio super-resolution (ASR) model is used the enhance the generated audio quality and recover the losses occurred due to rescaling. More details on ASR are given in the supplementary. We scale and invert this enhanced audio spectrogram using Griffin-Lim [35] algorithm to obtain the audio waveform of 10 seconds. The average infernce time for our method is at par with the current SOTA V2A model [32]. The inference time comparison is presented in Table 1.
@@ 4 Experiments
Table 1: V2A generation results comparing various methods over 3 metrics (FID, IS, VAAS). We have also compared them against the audio sampling rate (ASR), maximum frequency support (MaxF), spectrogram size used (Size), the length of generated audio (AL) in seconds, the number of video frames used for guidance (VF), and inference time (IT) in seconds. Our method, using 30 fps, sampled at highest frequency can generate the highest resolution spectrogram for a 10 seconds long audio. It also outperforms the existing V2A on FID, IS, and VAAS.
Method ASR MaxF Size AL (sec) VF FID $\downarrow$ IS $\uparrow$ VAAS $\uparrow$ IT (sec)
Im2Wav [45] $16KHz$ $8KHz$- $5$ $1-30fps$ $91.73$ $19.15$ $59.24$ $22$
Make-an-Audio [23] $16KHz$ $8KHz$ $80\times 624$ $9-10$ $1-4$ $83.72$ $22.14$ $64.57$ $17$
Diff-Foley [32] $16KHz$ $8KHz$ $256\times 128$ $8$ $4fps$ $76.92$ $37.92$ $78.19$ $1.9$
EgoSonics (ours) $22KHz$ $20KHz$ $512\times 512$ $10$ $30fps$ 41.14 54.41 92.77 $2.6$
@@
Table 2: Comparing the effect of using audio generated from different methods on improving video summarization task. It is interesting to note that, even the unsynchronized audios generated from our model have sufficient contextual information to outperform the existing methods.
@@
Im2Wav [45] $0.57$
Make-an-Audio [23] $0.68$
Diff-Foley [32] $0.71$
@@ 4.3 Ablation Study
Unlike previous methods using an average video embedding, or 1-4 fps for audio generation, we use 30 fps to extract the temporal information from the video. In this section, we analyze the impact of using 30 frames on AV synchronization. We compare the performance of SyncroNet at different fps including 4 fps used by the current SOTA Diff-Foley [32]. From Table 3, it can be seen that our method not only generates good quality audio spectrograms, but also achieves a VAAS score of 85. As we increase the fps to 15, the alignment score increases to 89, and then to 92.77 with 30 fps.
4.4 User Study
[…]
@@ 5 Conclusion
In this paper, we propose a novel audio-visual framework called EgoSonics. Much like how humans effortlessly connect sound with visual stimuli, EgoSonics has the ability to predict synchronized audio for muted Egocentric videos. Leveraging a streamlined architecture that incorporates generation capabilities of Stable Diffusion, EgoSonics holds promise for various applications. The proposed Syncronet can generate control signals from input video frames to provide pixel-level control to Stable Diffusion. Furthermore, we showcased our method’s superior performance over existing methods and illustrate how this approach can enhance downstream applications like video summarization.
@@
Although, EgoSonics can generate good quality audio for egocentric videos, there are a few limitations. One thing we observed in some of our generated samples is the misalignment due to the lack of rich visual information resulting from occlusions. Our approach is also limited by the amount of training data. For example, there are a very few samples of musical instruments in the Ego4D dataset, consequently, our model doesn’t perform very well on such videos. We believe that such challenges can be solved by training our model with millions of audio-video pairs.
A. User Study
We conducted an user study for a subjective evaluation of our synthesized audios. With randomly selected 15 participants from different backgrounds and asked them to do a 10 minutes survey using Google Forms. The survey presented 16 videos with audios to the participants and asked them to rate their realism on a scale of 1-5 (1 being ”Doesn’t Sound Real” and 2 being ”Sounds Real”). Participants were also asked if the audio is contextually relevant to the video, if the events in the video and the corresponding audio are synchronized. To verify that our method doesn’t overfit to one category of sounds, we also asked the participants to evaluate if the audio seems to be from a different category (e.g., a carpentry sound coming from a vacuum cleaner).
The following are the outcomes of this study:
  1. 1.
[…]
  2. 2.
[…]
  3. 3.
[…]
  4. 4.
    In more than $80\%$ of the cases, users did not find our audios to be belonging to another category, meaning that EgoSonics doesn’t overfit to any particular class and is able to generate all kinds of daily activity audios.